OpenStudy (anonymous):

Factor the trinomial. 3(a + b)2 + 5(a + b) − 2

OpenStudy (anonymous):

am I able to use the quadratic equation with these values??

hartnn (hartnn):

just for sake of simplicity let x=a+b so now u have quadratic equation: 3x^2+5x-2 now u can use quad eq. .....

hartnn (hartnn):

in the end u can resubstitute back x as x=a+b

OpenStudy (anonymous):

alright, so A+B= \[(-5\pm7)/6\] where do I go from here?

OpenStudy (anonymous):

my two answers lead to -2 and 1/3. How do I present my results in an answer form?

hartnn (hartnn):

if a is the root x-a is the factor make sense?

hartnn (hartnn):

so if x=1/3 factor will be x-1/3 or (3x-1) so (3a+3b-1) can u find other factor???

OpenStudy (anonymous):

I'm failing to understand the concept. I appreciate that you're not solving the problem for me, but what did you do to transform your x-1/3 to (3x-1)?

hartnn (hartnn):

x-1/3 = (3x-1)/3 right? since f(x)=0 was your quadratic equation, that /3 will not have any effect.

hartnn (hartnn):

the other factor would be x+2 (a+b+2)